OpenStudy (thebadyboy):

Which sets of numbers are closed under addition? Choose all answers that are correct. A. {0, 2, 5, 8} B. even integers C. rational numbers D. {0}

OpenStudy (thebadyboy):

@PyroYolka @Ilovecake

OpenStudy (mimi_x3):

k well closed under addition means that when you add any of the values stated in that set the sum will still be in that set For instance lets begin with A {0,2,5,8} 0+2=2 Since 2 is in the set we can still assume its closed under addition 2+5=7 Is 7 in the set of {0,2,5,8} Obviously notttttt sooo therefore we can conclude that this set is NOT closed under addition

OpenStudy (mimi_x3):

Now lets look at the second option All even integers Even integers can be represented as 2k where k is an integer Now lets add 2 integers 2l+2m where l and m are both integers 2l+2m=2(l+m) where l and m are integers Therefore we can conclude that when adding even integers it will always result in an even integer. Hence it is closed under addition

OpenStudy (mimi_x3):

@Thebadyboy Do you follow so far?

OpenStudy (thebadyboy):

yes

OpenStudy (mimi_x3):

Lets go to the third option which is rational numbers Rational numbers is a real number that can be written as a fraction Rational numbers can be represented as \(\frac{p}{q}\) where p and q are integers Now lets add 2 rational numbers \(\frac{a}{b}+\frac{c}{d}\) where a,b,c and d are integers \(\frac{a}{b}+\frac{c}{d}=\frac{ad+cb}{bd}\) Since a,b, c and d are integers we can conclude that \(\frac{ad+cb}{bd}\) is a rational number Hence it is closed under addition

OpenStudy (thebadyboy):

that makes sence

OpenStudy (mimi_x3):

Now the last option is pretty obvious The set only containing 0 0+0=0 Hence .... It is closed under addition

OpenStudy (thebadyboy):

thx

OpenStudy (mimi_x3):

not sure what grade you are in but if you didnt understand the second and third option I can explain it in simpler terms

OpenStudy (thebadyboy):

thx i understand

OpenStudy (thebadyboy):

Numbers p and q are positive integers. Which statements are always true? Choose all answers that are correct. A. p + q is a positive integer B. p – q is a negative integer C. p • q is a positive integer D. p/q is a positive integer

OpenStudy (thebadyboy):

can u answer this one plz

OpenStudy (mimi_x3):

Lol can u attempt to answer this ... ill guide u

OpenStudy (thebadyboy):

ok